# PKU-DyMVHumans: A Multi-View Video Benchmark for High-Fidelity Dynamic Human Modeling

### v1

@@ Abstract
High-quality human reconstruction and photo-realistic rendering of a dynamic scene is a long-standing problem in computer vision and graphics. Despite considerable efforts invested in developing various capture systems and reconstruction algorithms, recent advancements still struggle with loose or oversized clothing and overly complex poses. In part, this is due to the challenges of acquiring high-quality human datasets. To facilitate the development of these fields, in this paper, we present PKU-DyMVHumans, a versatile human-centric dataset for high-fidelity reconstruction and rendering of dynamic human scenarios from dense multi-view videos. It comprises 8.2 million frames captured by more than 56 synchronized cameras across diverse scenarios. These sequences comprise 32 human subjects across 45 different scenarios, each with a high-detailed appearance and realistic human motion. Inspired by recent advancements in neural radiance field (NeRF)-based scene representations, we carefully set up an off-the-shelf framework that is easy to provide those state-of-the-art NeRF-based implementations and benchmark on PKU-DyMVHumans dataset. It is paving the way for various applications like fine-grained foreground/background decomposition, high-quality human reconstruction and photo-realistic novel view synthesis of a dynamic scene. Extensive studies are performed on the benchmark, demonstrating new observations and challenges that emerge from using such high-fidelity dynamic data. The dataset is available at: https://pku-dymvhumans.github.io.
@@ 1 Introduction
In this work, we propose PKU-DyMVHumans, a versatile human-centric dataset that includes 32 dynamic humans performing various actions and wearing different clothing styles, as shown in Fig. 1. The dataset consists of approximately 8.2 million frames and aims to address the lack of large-scale and high-fidelity human performance datasets. Compared to existing benchmark datasets, PKU-DyMVHumans dataset offers appealing characteristics: 1) High-fidelity human performance: We construct professional multi-view system to capture humans in motion, which contains 56/60 synchronous cameras with 1080P and 4K resolution. 2) High-detailed appearance: It captures complex cloth deformation, and intricate texture details, such as delicate cloth wrinkles and vivid textures of classical headwear. 3) Complex human motion: It covers a wide range of special costume performances, artistic movements, and sports activities. 4) Real-human-scene interactions: These include human-object interactions, as well as challenging multi-person interactions and complex scene effects (e.g., lighting, shadows, and smoking).
@@ 2 Related Work
Dataset #Cam. #Subj. #Subj./img #Scene #Seq. #Frame Dynamic Interaction High-Detail Resolution
Human3.6M[11] 4 11 Single 15 839 3.6M 1000×1002
THUman2.0[49] 60 200 Single - - - 512×512
Hi4D[47] 8 40 Pair 20 100 11K 940×1280
DNA-Rendering[39] 60 500 Single 1,187 25,320 67.5M 4096×3000, 2448×2048
DynaCap[9] 50-101 4 Single 5 440 11.8M 1285×940
MultiHuman[54] 128 50 Multiple - - - -
THuman5.0[35] 32 10 Single 10 320 96K 4096×3000
HuMMan[2] 11 132 Single 20 4,466 278K 1920×1080
UltraStage[56] 32 100 Single 20 - 192K 7860×4320
Actors-HQ[10] 160 8 Single 21 2,560 39.8K 4096×3072
HUMBI[50] 107 772 Single 4 - 26M 1920×1080
NHR[45] 56/72 4 Single 3 320 47K 1024×768
AIST++[38] 9 30 Multiple 10 1,408 10.1M 1920×1080
ZJU_Mocap[31] 21 9 Single 6 2̃07 185K 1024×1024
THuman4.0[55] 24 3 Single 3 320 250K 1330×1150
ENeRF-Outdoor[18] 18 7 Multiple 4 72 86K 1920×1080
FreeMan[40] 8 40 Single 123 8,000 11.3M 1920×1080
PKU-DyMVHumans (Ours) 56/60 32 Multiple 45 2,668 8.2M 3840×2160, 1920×1080
@@
Table 1: Comparisons of multi-view human datasets. We compare the proposed dataset with previous human-centric multi-view datasets in terms of scale, attribute, and resolution. The gray color indicates the marker-based capture data with reliable human motion or pre-scanned humans reconstructed from depth sensors or stereo camera arrays. Compared to existing datasets, PKU-DyMVHumans features dynamic sequences with large scale, intricate details, complex motions, and interactions of multiple subjects.
@@
Human performance capture and datasets. Human performance capture, which involves capturing the full pose and non-rigid surface deformation of people wearing regular clothing in a space-time coherent 4D manner, has revolutionized the film and gaming industry in recent years [8]. The early existing methods learn from high-quality data obtained from optical marker-based capture systems [11, 49]. Other recent methods utilize depth sensor [35, 47, 39] or volumetric data [8, 9] to achieve reliable human motion and geometry. However, the sophisticated setup restricts their practical deployment. With the emergence of neural rendering techniques, rendering realistic humans directly from images has become a trend. Such a setting [1] usually requires the dataset equipped with high-quality dense view images [50, 45] or accurate annotations like human body keypoints and foreground segmentation [38, 31, 55, 40]. However, recent methods still often struggle to reconstruct models accurately when the clothes are loose, oversized, or when the poses are too complex. Differing with these efforts, we aim to provide a versatile human-centric dataset designed for high-fidelity reconstruction and rendering of dynamic human performances from dense multi-view videos. The unfold comparisons between PKU-DyMVHumans and the existing datasets are given in Tab. 1. When compared to the published benchmark datasets, PKU-DyMVHumans includes dynamic sequences with large scale, intricate details, complex motions, and interactions of multiple subjects at high resolutions.
@@
Free View synthesis for dynamic scenes. Photo-realistic rendering of dynamic scenes from a set of input images are necessary for many applications. For dynamic scene modeling, some methods [17, 16, 46] consider dynamic scenes as a 4D domain and add the time dimension to the input spatial coordinate. This approach enables the implementation of space-time radiance fields. In the field of human rendering, several approaches [22, 31, 44] utilize human priors to model human motions and achieve free-viewpoint rendering of dynamic human performance. HumanNeRF [44] demonstrated the ability to render realistic humans from monocular video sequences. While these methods produce impressive rendering results, their training process is time-consuming. To address the need for fast dynamic scene reconstruction, NeuS2 [42] integrates multi-resolution hash encoding into neural SDF and proposes an incremental training strategy with a global transformation prediction component. This approach leverages shared geometry and appearance information in two consecutive frames. In another line of work, Tensor4D [36] proposes a hierarchical tri-projection decomposition method to learn high-quality neural representation for dynamic scenes from sparse-view videos. This method models a 4D tensor using nine 2D feature planes, capturing spatio-temporal information in a compact and memory-efficient manner. However, they generally handle videos with short frames.
@@ 3.1 Data Capturing and Processing
Foreground object segmentation. Each frame extracted from the original video is passed to BGMv2 [19] or RVM [20] to generate the binary foreground object mask. This not only improves dense reconstruction but also contributes to the subsequent step of dynamic human reconstruction. In this context, we present a self-supervised learning framework towards delicate segmentation by leveraging the cross-view consistency of neural implicit surface representation from a sparse set of posed images (Sec. 4.3).
@@ 3.2 Dataset Statistics and Distribution
In the 1080P sequences category, the dataset consists of 36 scenes with a resolution of 1920×1080. The duration ranges from 10 to 487 seconds. The dance sequences involve complex and highly deformable clothing, sports sequences involve interactions between players and props (such as battledore, ball, and taekwondo plank), and kungfu sequences correspond to different choreography types with corresponding costumes and props (such as Broadsword, Nanquan, Nunchuck, Taichi, etc.). In the 4K Studio category, the dataset includes 8 scenes with a resolution of 3840×2160. The duration ranges from 10 to 231 seconds. This category covers both solo fashion shows and pair/group performances, as well as basic choreography dances and advanced dances originally choreographed by dancers.
@@ 3.3 Benchmark Pipeline
Width Ratio >0.4
1080_Dance_Dunhuang_Pair_f14f15
21.38 0.959 0.042 26.37 0.974 0.035 25.34 0.967 0.044
1080_Sport_Taekwondo1_Pair_m11c21
23.72 0.970 0.037 32.50 0.989 0.019 27.12 0.981 0.029
1080_Kungfu_Sword_Single_m13
23.24 0.978 0.023 31.10 0.986 0.019 29.39 0.985 0.019
1080_Kungfu_Spear_Single_m13
22.53 0.973 0.033 31.85 0.989 0.018 28.16 0.985 0.021
Width Ratio $\in(0.2,0.4)$
1080_Kungfu_Fan_Single_m12
25.72 0.981 0.024 33.16 0.988 0.021 30.56 0.985 0.021
1080_Kungfu_Basic_Pair_c24c25
25.48 0.979 0.024 31.96 0.989 0.017 30.99 0.986 0.019
4K_Studios_Dance_Single_f20
22.67 0.981 0.028 30.50 0.986 0.026 31.67 0.989 0.018
1080_Dance_Dunhuang_Single_f12
25.19 0.978 0.024 31.46 0.983 0.019 30.61 0.985 0.020
4K_Studios_Show_Single_f16
20.38 0.975 0.042 34.49 0.990 0.012 34.33 0.993 0.012
4K_Studios_Show_Pair_f18f19
22.80 0.976 0.031 34.24 0.993 0.016 32.23 0.992 0.014
Width Ratio <0.2
1080_Sport_Football_Single_m11
24.91 0.983 0.017 29.83 0.982 0.018 30.50 0.986 0.016
1080_Dance_Banquet_Single_c23
26.73 0.984 0.016 36.20 0.993 0.009 35.56 0.993 0.009
1080_Kungfu_Tongbeiquan_Single_m13
23.59 0.980 0.024 32.19 0.991 0.016 31.61 0.988 0.017
1080_Dance_Tibetan_Single_c22
26.26 0.984 0.020 33.69 0.991 0.016 34.28 0.990 0.014
@@
Table 2: Results of novel view synthesis on PKU-DyMVHumans dataset.
@@ 4 Experiments
Figure 6: The results of space and time novel view rendering under short frame sequences, with dense (NeuS2) and sparse (Tensor4D) camera views respectively.
@@ 4.2 Dynamic Human Modeling
PSNR $\uparrow$ / SSIM $\uparrow$ / LPIPS $\downarrow$
PSNR $\uparrow$ / SSIM $\uparrow$ / LPIPS $\downarrow$
PSNR $\uparrow$ / SSIM $\uparrow$ / LPIPS $\downarrow$
PSNR $\uparrow$ / SSIM $\uparrow$ / LPIPS $\downarrow$
Width Ratio >0.4
29.53 / 0.984 / 0.025
25.87 / 0.974 / 0.032
28.62 / 0.980 / 0.025
27.59 / 0.970 / 0.040
Width Ratio $\in(0.2,0.4)$
30.55 / 0.983 / 0.022
30.05 / 0.985 / 0.021
32.04 / 0.989 / 0.017
31.56 / 0.982 / 0.028
Width Ratio <0.2
29.39 / 0.984 / 0.022
30.52 / 0.987 / 0.018
32.93 / 0.989 / 0.015
28.89 / 0.986 / 0.022
Average
29.82 / 0.984 / 0.023
28.81 / 0.982 / 0.024
31.20 / 0.986 / 0.019
29.38 / 0.979 / 0.030
Implementation details. We apply the task of dynamic human reconstruction with tho different input settings: 1) Dynamic reconstruction under densely captured videos: In this setting, each sequence contains 10 to 500 frames with 48 or 52 camera views for training and 8 camera views for testing. Given the multi-view videos of a moving object and camera parameters for each view, we use Instant-NGP and NeuS2 to recover per-frame reconstruction and dynamic reconstruction, respectively. 2) Dynamic reconstruction using sparse and fixed cameras: We use 16 cameras focusing on the front face, with the number of training frames ranging from 10 to 20. For this setting, we employ NeuS2 and Tensor4D to represent topologically varying objects. For quantitative evaluation, we calculate and average the PSNR, LPIPS, and SSIM scores over all frames.
@@ 4.3 Neural Scene Decomposition
Neural scene decomposition aims to effectively separate the foreground and background without any annotations, with downstream applications in multi-view object segmentation [4] and beyond, such as 3D/4D human reconstruction and rendering. Under conditions of multi-camera inputs, the structural, textural and geometrical consistency among each view can be leveraged to achieve fine-grained object segmentation. To make better use of the above information, we present a self-supervised learning framework towards delicate segmentation by combining the neural surface representation power from multi-view observations via volume rendering of SDF. This representation captures the compositional nature of scenes and provides additional inherent information, thus improving 3D human reconstruction.
@@
We train our method for each individual sequence and present some example results in Fig. 7. Our method generates detailed foreground alpha, resulting in high-quality geometry surfaces. This is evidenced by the distinct cloth wrinkles and natural shading. The detailed procedures and more comparison against relevant methods of PKU-DyMVHumans are illustrated in the supplementary material. It is worth noting that we achieve these outcomes solely by decoupling the scene into two complementary neural scene representation modules. Applying more advanced designs in the future will certainly enhance the effects.
Width ratio of human box BGMv2 [19] Our method

### v3

@@ Abstract
High-quality human reconstruction and photo-realistic rendering of a dynamic scene is a long-standing problem in computer vision and graphics. Despite considerable efforts invested in developing various capture systems and reconstruction algorithms, recent advancements still struggle with loose or oversized clothing and overly complex poses. In part, this is due to the challenges of acquiring high-quality human datasets. To facilitate the development of these fields, in this paper, we present PKU-DyMVHumans, a versatile human-centric dataset for high-fidelity reconstruction and rendering of dynamic human scenarios from dense multi-view videos. It comprises 8.2 million frames captured by more than 56 synchronized cameras across diverse scenarios. These sequences comprise 32 human subjects across 45 different scenarios, each with a high-detailed appearance and realistic human motion. Inspired by recent advancements in neural radiance field (NeRF)-based scene representations, we carefully set up an off-the-shelf framework that is easy to provide those state-of-the-art NeRF-based implementations and benchmark on PKU-DyMVHumans dataset. It is paving the way for various applications like fine-grained foreground/background decomposition, high-quality human reconstruction and photo-realistic novel view synthesis of a dynamic scene. Extensive studies are performed on the benchmark, demonstrating new observations and challenges that emerge from using such high-fidelity dynamic data. The project page and data is available at: https://pku-dymvhumans.github.io.
@@ 1 Introduction
In this work, we propose PKU-DyMVHumans, a versatile human-centric dataset that includes 32 dynamic humans performing various actions and wearing different clothing styles, as shown in Fig. 1. The dataset consists of approximately 8.2 million frames and aims to address the lack of large-scale and high-fidelity human performance datasets. Compared to existing benchmark datasets, PKU-DyMVHumans dataset offers appealing characteristics: 1) High-fidelity human performance: We construct a professional multi-view system to capture humans in motion, which contains 56/60 synchronous cameras with 1080P or 4K resolution. 2) High-detailed appearance: It captures complex cloth deformation, and intricate texture details, such as delicate satin ribbon and vivid textures of classical headwear. 3) Complex human motion: It covers a wide range of special costume performances, artistic movements, and sports activities. 4) Human-object/scene interactions: These include human-object interactions, as well as challenging multi-person interactions and complex scene effects (e.g., lighting, shadows, and smoking).
@@ 2 Related Work
Table 1: Comparisons of multi-view human datasets. We compare the proposed dataset with previous human-centric multi-view datasets in terms of scale, attribute, and resolution. The gray color indicates the marker-based capture data with reliable human motion or pre-scanned humans reconstructed from depth sensors or stereo camera arrays. Differing from these efforts, PKU-DyMVHumans features high-fidelity dynamic sequences with high-detailed appearance, complex human motions, as well as challenging multi-person interactions.
Dataset #Cam. #Subj. #Scene #Seq. #Frame Dynamic Interaction High-detail Multi-person Resolution
Human3.6M[11] 4 11 15 839 3.6M 1000×1002
THUman2.0[48] 60 200 - - - 512×512
Hi4D[46] 8 40 20 100 11K 940×1280
DNA-Rendering[38] 60 500 1,187 25,320 67.5M 4096×3000, 2448×2048
DynaCap[9] 50-101 4 5 440 11.8M 1285×940
MultiHuman[54] 128 50 - - - -
THuman5.0[34] 32 10 10 320 96K 4096×3000
HuMMan[2] 11 132 20 4,466 278K 1920×1080
UltraStage[56] 32 100 20 - 192K 7860×4320
Actors-HQ[10] 160 8 21 2,560 39.8K 4096×3072
HUMBI[49] 107 772 4 - 26M 1920×1080
NHR[44] 56/72 4 3 320 47K 1024×768
AIST++[37] 9 30 10 1,408 10.1M 1920×1080
ZJU_Mocap[30] 21 9 6 2̃07 185K 1024×1024
THuman4.0[55] 24 3 3 320 250K 1330×1150
ENeRF-Outdoor[18] 18 7 4 72 86K 1920×1080
FreeMan[39] 8 40 123 8,000 11.3M 1920×1080
PKU-DyMVHumans (Ours) 56/60 32 45 2,668 8.2M 3840×2160, 1920×1080
@@
Human performance capture and datasets. Human performance capture, which involves capturing the full pose and non-rigid surface deformation of people wearing regular clothing in a space-time coherent 4D manner, has revolutionized the film and gaming industry in recent years [8]. The early existing methods learn from high-quality data obtained from optical marker-based capture systems [11, 48]. Other recent methods utilize depth sensor [34, 46, 38] or volumetric data [8, 9] to achieve reliable human motion and geometry. However, the sophisticated setup restricts their practical deployment. With the emergence of neural rendering techniques, rendering realistic humans directly from images has become a trend. Such a setting [1] usually requires the dataset equipped with high-quality dense view images [49, 44] or accurate annotations like human body keypoints and foreground segmentation [37, 30, 55, 39]. However, recent methods still often struggle to reconstruct models accurately when the clothes are loose, oversized, or when the poses are too complex. Differing from these efforts, we aim to provide a versatile human-centric dataset designed for high-fidelity reconstruction and rendering of dynamic human performances from dense multi-view videos. The unfold comparisons between PKU-DyMVHumans and the existing datasets are given in Tab. 1. When compared to the published benchmark datasets, PKU-DyMVHumans features high-detailed appearance, complex motion, as well as challenging human-object interactions, multi-person interactions and complex scene effects (e.g., lighting, shadows, and smoking).
@@
Free view synthesis for dynamic scenes. Photo-realistic rendering of dynamic scenes from a set of input images is necessary for many applications. For dynamic scene modeling, some methods [17, 16, 45] consider dynamic scenes as a 4D domain and add the time dimension to the input spatial coordinate. This approach enables the implementation of space-time radiance fields. In the field of human rendering, several approaches [21, 30, 43] utilize human priors to model human motions and achieve free-viewpoint rendering of dynamic human. HumanNeRF [43] demonstrates the ability to render realistic humans from monocular video sequences. While these methods produce impressive rendering results, their training process is time-consuming. To address the need for fast dynamic scene reconstruction, NeuS2 [41] integrates multi-resolution hash encoding into SDF and proposes an incremental training strategy with a global transformation prediction component. This approach leverages shared geometry and appearance information in two consecutive frames. Besides, Tensor4D [35] proposes a hierarchical tri-projection decomposition method to learn high-quality dynamic scenes representation from sparse-view videos. This method models a 4D tensor using nine 2D feature planes, capturing spatio-temporal information in a compact and memory-efficient manner. However, they generally handle videos with short frames.
@@ 3.1 Data Capturing and Processing
Foreground object segmentation. Each frame extracted from the original video is passed to BGMv2 [19] to generate the binary foreground object mask. This not only improves dense reconstruction but also contributes to the subsequent step of dynamic human reconstruction. In this context, our contemporary work Surface-SOS [53] aims for delicate segmentation by leveraging the cross-view consistency of neural implicit surface representation from a sparse set of posed images (Sec. 4.3).
@@ 3.2 Dataset Statistics and Distribution
In the 1080P sequences category, the dataset consists of 36 scenes with a resolution of 1920×1080. The duration ranges from 10 to 487 seconds. The dance sequences involve complex and highly deformable clothing, sports sequences involve interactions between players and props (such as battledore, ball, and taekwondo plank), and kungfu sequences correspond to different choreography types with corresponding costumes and props (such as Broadsword, Nanquan, Nunchuck, Taichi, etc.). In the 4K Studio category, the dataset includes 8 scenes with a resolution of 3840×2160. The duration of each sequence ranges from 10 to 231 seconds. This category covers both solo fashion shows and pair/group performances, as well as basic choreography dances and advanced dances originally choreographed by dancers.
@@ 3.3 Benchmark Pipeline
Table 2: Quantitative performance of novel view synthesis on the proposed PKU-DyMVHumans dataset.
@@
Width Ratio >0.4 1080_Dance_Dunhuang_Pair_f14f15 21.38 0.959 0.042 26.37 0.974 0.035 25.34 0.967 0.044
1080_Sport_Taekwondo1_Pair_m11c21 23.72 0.970 0.037 32.50 0.989 0.019 27.12 0.981 0.029
1080_Kungfu_Sword_Single_m13 23.24 0.978 0.023 31.10 0.986 0.019 29.39 0.985 0.019
1080_Kungfu_Spear_Single_m13 22.53 0.973 0.033 31.85 0.989 0.018 28.16 0.985 0.021
Width Ratio $\in(0.2,0.4)$ 1080_Kungfu_Fan_Single_m12 25.72 0.981 0.024 33.16 0.988 0.021 30.56 0.985 0.021
1080_Kungfu_Basic_Pair_c24c25 25.48 0.979 0.024 31.96 0.989 0.017 30.99 0.986 0.019
4K_Studios_Dance_Single_f20 22.67 0.981 0.028 30.50 0.986 0.026 31.67 0.989 0.018
1080_Dance_Dunhuang_Single_f12 25.19 0.978 0.024 31.46 0.983 0.019 30.61 0.985 0.020
4K_Studios_Show_Single_f16 20.38 0.975 0.042 34.49 0.990 0.012 34.33 0.993 0.012
4K_Studios_Show_Pair_f18f19 22.80 0.976 0.031 34.24 0.993 0.016 32.23 0.992 0.014
Width Ratio <0.2 1080_Sport_Football_Single_m11 24.91 0.983 0.017 29.83 0.982 0.018 30.50 0.986 0.016
1080_Dance_Banquet_Single_c23 26.73 0.984 0.016 36.20 0.993 0.009 35.56 0.993 0.009
1080_Kungfu_Tongbeiquan_Single_m13 23.59 0.980 0.024 32.19 0.991 0.016 31.61 0.988 0.017
1080_Dance_Tibetan_Single_c22 26.26 0.984 0.020 33.69 0.991 0.016 34.28 0.990 0.014
@@ 4 Experiments
Figure 6: The results of space and time novel view rendering under short frame sequences, with dense (NeuS2 [41]) and sparse (Tensor4D [35]) camera views respectively.
@@ 4.2 Dynamic Human Modeling
PSNR $\uparrow$ / SSIM $\uparrow$ / LPIPS $\downarrow$ PSNR $\uparrow$ / SSIM $\uparrow$ / LPIPS $\downarrow$ PSNR $\uparrow$ / SSIM $\uparrow$ / LPIPS $\downarrow$ PSNR $\uparrow$ / SSIM $\uparrow$ / LPIPS $\downarrow$
Width Ratio >0.4 29.53 / 0.984 / 0.025 25.87 / 0.974 / 0.032 28.62 / 0.980 / 0.025 27.59 / 0.970 / 0.040
Width Ratio $\in(0.2,0.4)$ 30.55 / 0.983 / 0.022 30.05 / 0.985 / 0.021 32.04 / 0.989 / 0.017 31.56 / 0.982 / 0.028
Width Ratio <0.2 29.39 / 0.984 / 0.022 30.52 / 0.987 / 0.018 32.93 / 0.989 / 0.015 28.89 / 0.986 / 0.022
Average 29.82 / 0.984 / 0.023 28.81 / 0.982 / 0.024 31.20 / 0.986 / 0.019 29.38 / 0.979 / 0.030
Implementation details. We apply the task of dynamic human reconstruction with two different input settings: 1) Dynamic reconstruction under densely captured videos: In this setting, each sequence contains 10 to 500 frames with 48 or 52 camera views for training and 8 camera views for testing. Given the multi-view videos of a moving object and camera parameters for each view, we use Instant-NGP and NeuS2 to recover per-frame reconstruction and dynamic reconstruction, respectively. 2) Dynamic reconstruction using sparse and fixed cameras: We use 16 cameras focusing on the front face, with the number of training frames ranging from 10 to 20. For this setting, we employ NeuS2 and Tensor4D to represent topologically varying objects. For quantitative evaluation, we calculate and average the PSNR, LPIPS, and SSIM scores over all frames.
@@ 4.3 Neural Scene Decomposition
Neural scene decomposition aims to effectively separate the foreground and background without any annotations, with downstream applications in multi-view object segmentation [4] and beyond, such as 3D/4D human reconstruction and rendering. Under conditions of multi-camera inputs, the structural, textural and geometrical consistency among each view can be leveraged to achieve fine-grained object segmentation. Surface-SOS [53] addresses this information by combining the neural surface representation from multi-view observations via volume rendering of SDF. It captures the compositional nature of scenes and provides additional inherent information, improving 3D human reconstruction.
@@
We train Surface-SOS [53] for each individual sequence and present some example results in Fig. 7. It shows that Surface-SOS can generates detailed foreground alpha, resulting in high-quality geometry surfaces. This is evidenced by the distinct cloth wrinkles and natural shading. The quantitative results in Tab. 4 also highlights the improvement of the Surface-SOS refinement. The detailed procedures and more comparison result against relevant methods of PKU-DyMVHumans are further illustrated in the supplementary material. It shows that when dealing with scenes involving human-object interactions, multi-person interactions, as well as challenging complex scene effects, it faces even greater challenges.
Width ratio of human box BGMv2 [19] Surface-SOS [53]